A child who weighs 325 N stands 5 m from the right end of a 12 m plank. A 150 N load has its center of gravity 4 m from the same end. The plank weighs 110 N. If the plank is supported at its ends, find the force exerted by each support. Show solutions.


1
Expert's answer
2021-05-21T10:43:33-0400

"110\cdot 6+325\cdot 7+150\cdot 8=F_2\cdot 12,\\
F_2=345~N,"

"150\cdot 4+325\cdot 5+110\cdot 6=F_1\cdot 12,\\
F_1=240~N."